# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

### v1

@@ 2.2.1. Infinite-dimensional Wishart process
$\int_{0}^{t}\|e^{s\mathbb{A}}\sqrt{Q}\|^{2}_{2}\,\mathop{}\!\textnormal{d}s<\infty\,$,
@@
$\mathop{}\!\textnormal{d}\mathcal{Y}_{t}$ $=nQ\,\mathop{}\!\textnormal{d}t+\mathcal{Y}_{t}\mathbb{A}\mathop{}\!\textnormal{d}t+\mathbb{A}^{*}\mathcal{Y}_{t}\,\mathop{}\!\textnormal{d}t+\sqrt{\mathcal{Y}}_{t}\,\mathop{}\!\textnormal{d}B_{t}\sqrt{Q}+\sqrt{Q}\mathop{}\!\textnormal{d}B_{t}^{*}\sqrt{\mathcal{Y}}_{t}\,,\quad t\geq 0\,$,
@@
where $(B_{t})_{t\geq 0}$ is an $\mathcal{L}_{2}(H)$-cylindrical Brownian motion. As shown in [14, Theorem 2.1], equation (2.2.1) admits a probabilistically and analytically weak solution. In fact, the existence result in [14, Theorem 2.1] holds more generally when $\mathbb{A}$ is an unbounded operator generating a strongly-continuous semigroup.
Furthermore, [14, Proposition 4.7 and Corollary 4.11] provide a characterization of finite-rank $\mathcal{S}_{1}^{+}(H)$-valued Wishart processes. In this paper, however, we restrict our attention to the case of a bounded operator $\mathbb{A}$, since under this assumption equation (2.2.1) can naturally serve as a model for the instantaneous covariance process as we argue next.
@@ 2.2.2. The infinite-dimensional Wishart stochastic volatility model
Let $n\in\mathbb{N}$, let $y\in\mathcal{S}_{1}^{+}(H)$ be of rank at most $n$, and let $(W_{t})_{t\geq 0}$ be an $H$-cylindrical Brownian motion. Let $x,\Upsilon\in H$ and $D\in\mathcal{S}_{1}^{+}(H)$, and let $(X_{t})_{t\geq 0}$ be given as the solution to the HJMM–SPDE (1.1).
@@
$\begin{cases}\mathop{}\!\textnormal{d}X_{t}=\big{(}\mathcal{A}X_{t}+D^{1/2}\mathcal{Y}_{t}D^{1/2}\Upsilon\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}\mathcal{Y}_{t}^{1/2}\,\mathop{}\!\textnormal{d}W_{t},\\[4.0pt] \mathop{}\!\textnormal{d}\mathcal{Y}_{t}=\big{(}nQ+\mathcal{Y}_{t}\mathbb{A}+\mathbb{A}^{*}\mathcal{Y}_{t}\big{)}\,\mathop{}\!\textnormal{d}t+\sqrt{\mathcal{Y}_{t}}\,\mathop{}\!\textnormal{d}B_{t}\sqrt{Q}+\sqrt{Q}\,\mathop{}\!\textnormal{d}B_{t}^{*}\sqrt{\mathcal{Y}_{t}},\quad t\geq 0,\end{cases}$ (2.10)
where $(W_{t})_{t\geq 0}$ and $(B_{t})_{t\geq 0}$ are independent. The existence of a (stochastically) weak solution to (2.10) in $H\times\mathcal{S}_{1}^{+}(H)$ follows from [16, Lemma 2.8] and [14, Theorem 2.1].
Following [16], and motivated by applications, we include the linear term $D^{1/2}\mathcal{Y}_{t}D^{1/2}\Upsilon$ in the drift of $(X_{t})_{t\geq 0}$. This addition does not affect well-posedness or the affine property; see, for example, the proof of [16, Lemma 2.8]. We refer to (2.10) as the infinite-dimensional Wishart stochastic volatility model.
@@ 2.2.3. Riccati equations associated with the infinite-dimensional Wishart stochastic volatility model
Let $n\in\mathbb{N}$, $Q,D\in\mathcal{S}_{1}^{+}(H)$, $\Upsilon\in H$, and $\mathbb{A}\in\mathcal{L}(H)$. Define $F\colon\mathcal{S}_{1}(H)\to\mathbb{R}$ and $R\colon H\times\mathcal{S}_{1}(H)\to\mathcal{L}_{1}(H)$ by
@@
As in the affine pure-jump case, the functions $F$ and $R$ give rise to a system of operator-valued Riccati-type equations. Since the operators $(\mathcal{Y}_{t})_{t\geq 0}$ of the Wishart process are, by construction, of finite rank, we adopt a numerically convenient simplification: we approximate the model by a matrix-valued Wishart process whose rank remains constant over time. This reduces the infinite-dimensional problem (at least in the covariance dimensions) to a more tractable finite-dimensional one while preserving the essential structure of the model.
Let $(H_{n})_{n\in\mathbb{N}}$ denote an increasing sequence of finite-dimensional subspaces such that $\overline{\bigcup_{n}H_{n}}=H$. Let $\{e_{1},\ldots,e_{n}\}$ be an orthonormal basis of $H_{n}$ and define the orthogonal projection $\Pi_{n}:H\to H_{n}$. We denote by $\mathcal{L}_{n}:=\mathcal{L}(H_{n})\cong\mathbb{R}^{n\times n}$ the space of bounded linear operators on $H_{n}$ and by $\mathcal{S}_{n}$ the space of real symmetric $n\times n$ matrices. Define the finite-rank counterparts $F_{n}\colon\mathcal{S}_{n}\to\mathbb{R}$ and $R_{n}\colon H\times\mathcal{S}_{n}\to\mathcal{L}(H_{n})$ of the nonlinear maps $F$ and $R$ introduced in (2.11) and (2.12) via
@@
Let $F_{n}$ and $R_{n}$ be as in (2.13). Then, for $u=(u_{1},u_{2})\in H\times\mathcal{S}_{n}$ and $T\geq 0$, the finite-rank Riccati system is given by the system of ordinary differential equations
@@
$q_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}q_{1}(s,u)\,\mathrm{d}s,\quad$ $q_{1}(0,u)=u_{1},$ (2.14b)
@@
We call $(P_{n},q_{1},q_{2,n})$ a solution to (2.14) if $P_{n}\in C^{1}([0,T],\mathbb{R})$, $q_{1}\in C^{1}([0,T],H)$, and $q_{2,n}\in C^{1}([0,T],\mathcal{S}_{n})$ solve (2.14a)–(2.14c).
@@
Denote by $\mathcal{S}_{n}^{\mathbb{C}}=\mathcal{S}_{n}\oplus\mathrm{i}\mathcal{S}^{n}$. Let $F_{n}$ be the analytic extension to $\mathcal{S}_{n}^{\mathbb{C}}$ of the function $F_{n}$ defined in (2.13) and $R_{n}$ be the analytic extension to $\mathbb{C}^{n}\times\mathcal{S}_{n}^{\mathbb{C}}$ of the function $R_{n}$ defined in (2.13). Consider the following system of ordinary differential equations
@@
$\,\zeta_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\zeta_{1}(s,u)\mathop{}\!\textnormal{d}s,\quad 0<t\leq T,\quad\zeta_{1}(0,u)=u_{1}\,$, (2.15b)
@@
For $u=(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{S}_{n}^{\mathbb{C}}$ and $T\geq 0$, we say that $(\eta_{n}(\cdot,u),\zeta_{1}(\cdot,u),\zeta_{2,n}(\cdot,u))\colon[0,T]\rightarrow\mathbb{R}\times H^{\mathbb{C}}\times\mathcal{S}_{n}^{\mathbb{C}}$ is a mild solution to (2.15a)-(2.15c) if $\eta_{n}(\cdot,u)\in C^{1}([0,T],\mathbb{R})$, $\zeta_{1}(\cdot,u)\in C^{1}([0,T],H^{\mathbb{C}})$ and $\zeta_{2,n}(\cdot,u)\in C^{1}([0,T],\mathcal{S}_{n}^{\mathbb{C}})$ satisfies (2.15a)-(2.15c). Then we call $(\eta_{n},\zeta_{1},\zeta_{2,n})$ a solution (up to time $T$ and with starting point $u$) of a complex Riccati system.
Next, we state an existence result for the complex Riccati system (2.15). The proof follows the same lines as Proposition 5.5; see also [32, Proposition 5.1] for the case $D=S_{d}^{+}$, $d\geq 2$.
@@
Let $T>0$ and $n\in\mathbb{N}$. Let $Q,D\in\mathcal{S}_{1}^{+}(H)$, $\Upsilon\in H$, and $\mathbb{A}\in\mathcal{L}(H)$. Let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup $(S_{t})_{t\geq 0}$ on $H$, and $(\mathcal{A}^{*},\operatorname{dom}(\mathcal{A}^{*}))$ its adjoint. Let $u=(u_{1},u_{2})\in\mathbb{C}^{n}\times(\mathcal{S}_{n}^{+})^{\mathbb{C}}$ and suppose that the real Riccati system (2.14a)–(2.14c) with initial data $\Re(u)=(\Re(u_{1}),\Re(u_{2}))$ admits a (unique) mild solution $(P_{n},q_{1,n},q_{2,n})$ on $[0,T]$, as guaranteed by Proposition 5.6. Then the complex Riccati system (2.15a)–(2.15c) with initial data $u$ admits a (necessarily unique) mild solution $(\eta_{n},\zeta_{1,n},\zeta_{2,n})$ on $[0,T]$.
@@
To keep the presentation concise and highlight the numerical methodology, we approximate the Laplace transform of the infinite-dimensional Wishart stochastic volatility model directly. Let $(X_{t},\mathcal{Y}_{t})_{t\geq 0}$ be as in (2.10), and let $(\eta_{n},\zeta_{1},\zeta_{2,n})$ be the solution to (2.15) with initial condition $(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{S}_{n}^{\mathbb{C}}$, whose existence is guaranteed by Proposition 2.6.
@@
$\mathbb{E}\!\left[e^{\langle X_{T},u_{1}\rangle_{H^{\mathbb{C}}}-\operatorname{Tr}\big{(}\mathcal{Y}_{T}\,\iota_{n}u_{2}\iota_{n}^{*}\big{)}}\,\middle|\,\mathcal{F}_{t}\right]$
@@ 3. Pricing options on forwards in infinite-dimensional stochastic volatility models
Suppose $w\colon\mathbb{R}^{+}\rightarrow[1,\infty)$ is a non-decreasing function such that $w(0)=1$ and $\int_{0}^{\infty}w^{-1}(x)\,\mathop{}\!\textnormal{d}x<\infty$. Let $H_{w}$ be the space of absolutely continuous functions $f:\mathbb{R}^{+}\rightarrow\mathbb{R}$ such that
$\|f\|_{w}^{2}:=f(0)^{2}+\int_{0}^{\infty}w(x)|f^{\prime}(x)|^{2}\,\mathop{}\!\textnormal{d}x<\infty\,$,
@@
$\langle f,g\rangle_{w}:=f(0)g(0)+\int_{0}^{\infty}w(x)f^{\prime}(x)g^{\prime}(x)\,\mathop{}\!\textnormal{d}x.$
@@
The left-shift semigroup is a strongly continuous semigroup on $H_{w}$ with infinitesimal generator $\mathcal{A}$ being the operator of differentiation (in time-to-maturity coordinate). Moreover, the point evaluation $\delta_{x}(h)=u_{x}$ is a continuous linear functional on $H_{w}$, for all $x\in\mathbb{R}^{+}$. It can be expressed as (see [25, Lemma 5.3.1])
@@
where for $x\in\mathbb{R}^{+}$, $u_{x}:\mathbb{R}^{+}\rightarrow\mathbb{R}$, $y\mapsto 1+\int_{0}^{x\wedge y}w^{-1}(z)\,\mathop{}\!\textnormal{d}z$.
@@
$\mathop{}\!\textnormal{d}X_{t}=\big{(}\mathcal{A}X_{t}+C_{t}\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}Y_{t}^{1/2}\,\mathop{}\!\textnormal{d}W_{t},\qquad X_{0}=x\in H_{w},$ (3.2)
where $\mathcal{A}=\partial_{x}$ is the derivative (shift) operator, $C$ is an integrable adapted $H_{w}$–valued process, $D\in\mathcal{L}(H_{w})$ is self-adjoint and positive, $(W_{t})_{t\geq 0}$ is a cylindrical Brownian motion, and $(Y_{t})_{t\geq 0}$ is either $(\mathcal{Y}_{t})_{t\geq 0}$ from (2.2.1) or $({\sf Y}_{t})_{t\geq 0}$ from (2.2). On a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{P})$, we model the (arbitrage-free) forward curve via
@@
$\mathbb{E}\left[\exp\!\Big{\{}-\int_{0}^{T}\langle\xi_{s},\mathop{}\!\textnormal{d}W_{s}\rangle_{w}-\tfrac{1}{2}\int_{0}^{T}\|\xi_{s}\|_{w}^{2}\,\mathop{}\!\textnormal{d}s\Big{\}}\right]=1.$ (3.4)
By Girsanov’s theorem for cylindrical Brownian motion (e.g., [20, Thm. 13]),
$\widetilde{W}_{t}:=W_{t}-\int_{0}^{t}\xi_{s}\,\mathop{}\!\textnormal{d}s$
@@
$\left.\frac{\mathop{}\!\textnormal{d}\mathbb{Q}}{\mathop{}\!\textnormal{d}\mathbb{P}}\right|_{\mathcal{F}_{t}}=\exp\Big{\{}-\int_{0}^{t}\langle\xi_{s},\mathop{}\!\textnormal{d}W_{s}\rangle_{w}-\tfrac{1}{2}\int_{0}^{t}\|\xi_{s}\|_{w}^{2}\,\mathop{}\!\textnormal{d}s\Big{\}},\qquad 0\leq t\leq T.$
@@
$\mathop{}\!\textnormal{d}X_{t}=\big{(}\mathcal{A}X_{t}+C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}Y_{t}^{1/2}\,\mathop{}\!\textnormal{d}\widetilde{W}_{t},\qquad 0\leq t\leq T.$ (3.5)
@@
. Let $\xi$ satisfy (3.4). Let $\mathcal{A}$, $C$, $D$, and $(Y_{t})_{t\geq 0}$ be as above, and let $S(t)$ be the left–shift semigroup generated by $\mathcal{A}$. Suppose there exists $h_{0}\in H_{w}$ such that, for all $0\leq t\leq\tau\leq T$,
$C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}=-\tfrac{1}{2}\,D^{1/2}Y_{t}D^{1/2}\,u_{\tau-t}\qquad\text{with }u_{x}:=S^{*}(x)h_{0}.$ (3.6)

### v2

@@ 2.2.1. Infinite-dimensional Wishart process
$\int_{0}^{t}\|\mathop{}\!\mathrm{e}^{s\mathbb{A}}\sqrt{Q}\|^{2}_{2}\,\mathop{}\!\mathrm{d}s<\infty\,$,
@@
$\mathop{}\!\mathrm{d}\mathcal{Y}_{t}$ $=nQ\,\mathop{}\!\mathrm{d}t+\mathcal{Y}_{t}\mathbb{A}\mathop{}\!\mathrm{d}t+\mathbb{A}^{*}\mathcal{Y}_{t}\,\mathop{}\!\mathrm{d}t+\sqrt{\mathcal{Y}}_{t}\,\mathop{}\!\mathrm{d}B_{t}\sqrt{Q}+\sqrt{Q}\mathop{}\!\mathrm{d}B_{t}^{*}\sqrt{\mathcal{Y}}_{t}\,,\quad t\geq 0\,$,
@@
where $(B_{t})_{t\geq 0}$ is an $\mathcal{L}_{2}(H)$-cylindrical Brownian motion. As shown in [13, Theorem 2.1], equation (2.2.1) admits a probabilistically and analytically weak solution. In fact, the existence result in [13, Theorem 2.1] holds more generally when $\mathbb{A}$ is an unbounded operator generating a strongly-continuous semigroup.
Furthermore, [13, Proposition 4.7 and Corollary 4.11] provide a characterization of finite-rank $\mathcal{S}_{1}^{+}(H)$-valued Wishart processes. In this paper, however, we restrict our attention to the case of a bounded operator $\mathbb{A}$, since under this assumption equation (2.2.1) can naturally serve as a model for the instantaneous covariance process as we argue next.
@@ 2.2.2. The infinite-dimensional Wishart stochastic volatility model
Let $n\in\mathbb{N}$, let $y\in\mathcal{S}_{1}^{+}(H)$ be of rank at most $n$, and let $(W_{t})_{t\geq 0}$ be an $H$-cylindrical Brownian motion. Let $x,\Upsilon\in H$ and $D\in\mathcal{S}_{1}^{+}(H)$.
@@
$\begin{cases}\mathop{}\!\mathrm{d}X_{t}=\big(\mathcal{A}X_{t}+D^{1/2}\mathcal{Y}_{t}D^{1/2}\Upsilon\big)\,\mathop{}\!\mathrm{d}t+D^{1/2}\mathcal{Y}_{t}^{1/2}\,\mathop{}\!\mathrm{d}W_{t},\\[4.0pt] \mathop{}\!\mathrm{d}\mathcal{Y}_{t}=\big(nQ+\mathcal{Y}_{t}\mathbb{A}+\mathbb{A}^{*}\mathcal{Y}_{t}\big)\,\mathop{}\!\mathrm{d}t+\sqrt{\mathcal{Y}_{t}}\,\mathop{}\!\mathrm{d}B_{t}\sqrt{Q}+\sqrt{Q}\,\mathop{}\!\mathrm{d}B_{t}^{*}\sqrt{\mathcal{Y}_{t}},\quad t\geq 0,\end{cases}$ (2.10)
where $(W_{t})_{t\geq 0}$ and $(B_{t})_{t\geq 0}$ are independent. The existence of a (stochastically) weak solution to (2.10) in $H\times\mathcal{S}_{1}^{+}(H)$ follows from [15, Lemma 2.8] and [13, Theorem 2.1]. Similarly to the pure-jump stochastic volatility model, we include the linear term $D^{1/2}\mathcal{Y}_{t}D^{1/2}\Upsilon$ in the drift of $(X_{t})_{t\geq 0}$, as motivated by the no-arbitrage drift condition; see Section 3.
We refer to (2.10) as the infinite-dimensional Wishart stochastic volatility model.
@@ 2.2.3. Riccati equations associated with the infinite-dimensional Wishart stochastic volatility model
Let $n\in\mathbb{N}$, $Q,D\in\mathcal{S}_{1}^{+}(H)$, $\Upsilon\in H$, and $\mathbb{A}\in\mathcal{L}(H)$. Define $F\colon\mathcal{L}_{1}(H)\to\mathbb{R}$ and $R\colon H\times\mathcal{L}_{1}(H)\to\mathcal{L}_{1}(H)$ by
@@
As in the affine pure-jump case, the functions $F$ and $R$ give rise to a system of operator-valued Riccati-type equations. Because of the mixed term $D^{1/2}h\otimes D^{1/2}\Upsilon$, the Riccati vector field $R(h,\cdot)$ does not preserve self-adjointness in general. Accordingly, the Riccati variable is matrix-valued, even though the Wishart state process $(\mathcal{Y}_{t})_{t\geq 0}$ itself remains in the cone of self-adjoint positive operators. Since the operators $(\mathcal{Y}_{t})_{t\geq 0}$ of the Wishart process are finite rank in the present setup once the initial condition has rank at most $n$, we adopt a numerically convenient simplification: we approximate the model by a matrix-valued Wishart process whose rank remains constant over time. This reduces the infinite-dimensional problem in the covariance component to a more tractable finite-dimensional one while preserving the maturity-space structure of the forward curve model.
Let $(H_{n})_{n\in\mathbb{N}}$ denote an increasing sequence of finite-dimensional subspaces such that $\overline{\bigcup_{n}H_{n}}=H$. Let $\{e_{1},\ldots,e_{n}\}$ be an orthonormal basis of $H_{n}$ and define the orthogonal projection $\Pi_{n}:H\to H_{n}$. We denote by $\mathcal{L}_{n}:=\mathcal{L}(H_{n})\cong\mathbb{R}^{n\times n}$ the space of bounded linear operators on $H_{n}$ and by $\mathcal{S}_{n}$ the space of real symmetric $n\times n$ matrices. Define the finite-rank counterparts $F_{n}\colon\mathcal{L}_{n}\to\mathbb{R}$ and $R_{n}\colon H\times\mathcal{L}_{n}\to\mathcal{L}_{n}$ of the nonlinear maps $F$ and $R$ introduced in (2.11) and (2.12) via
@@
Let $F_{n}$ and $R_{n}$ be as in (2.13). Then, for $u=(u_{1},u_{2})\in H\times\mathcal{L}_{n}$ and $T\geq 0$, the finite-rank Riccati system is given by the system of ordinary differential equations
@@
$q_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}q_{1}(s,u)\,\mathop{}\!\mathrm{d}s,\quad$ $q_{1}(0,u)=u_{1},$ (2.14b)
@@
We call $(P_{n},q_{1},q_{2,n})$ a solution to (2.14) if $P_{n}\in C^{1}([0,T],\mathbb{R})$, $q_{1}\in C^{1}([0,T],H)$, and $q_{2,n}\in C^{1}([0,T],\mathcal{L}_{n})$ solve (2.14a)–(2.14c).
@@
Denote by $\mathcal{L}_{n}^{\mathbb{C}}=\mathcal{L}_{n}\oplus\mathrm{i}\mathcal{L}_{n}$. Let $F_{n}$ be the analytic extension to $\mathcal{L}_{n}^{\mathbb{C}}$ of the function $F_{n}$ defined in (2.13) and $R_{n}$ be the analytic extension to $H^{\mathbb{C}}\times\mathcal{L}_{n}^{\mathbb{C}}$ of the function $R_{n}$ defined in (2.13). Consider the following system of ordinary differential equations
@@
$\,\zeta_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\zeta_{1}(s,u)\mathop{}\!\mathrm{d}s,\quad 0<t\leq T,\quad\zeta_{1}(0,u)=u_{1}\,$, (2.15b)
@@
For $u=(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{L}_{n}^{\mathbb{C}}$ and $T\geq 0$, we say that $(\eta_{n}(\cdot,u),\zeta_{1}(\cdot,u),\zeta_{2,n}(\cdot,u))\colon[0,T]\rightarrow\mathbb{R}\times H^{\mathbb{C}}\times\mathcal{L}_{n}^{\mathbb{C}}$ is a mild solution to (2.15a)-(2.15c) if $\eta_{n}(\cdot,u)\in C^{1}([0,T],\mathbb{R})$, $\zeta_{1}(\cdot,u)\in C^{1}([0,T],H^{\mathbb{C}})$ and $\zeta_{2,n}(\cdot,u)\in C^{1}([0,T],\mathcal{L}_{n}^{\mathbb{C}})$ satisfies (2.15a)-(2.15c). Then we call $(\eta_{n},\zeta_{1},\zeta_{2,n})$ a solution (up to time $T$ and with starting point $u$) of the complex Riccati system.
Next, we state an existence result for the complex Riccati system (2.15). The proof follows the same lines as Proposition 5.5; see also [31, Proposition 5.1] for the case $D=S_{d}^{+}$, $d\geq 2$.
@@
Let $T>0$ and $n\in\mathbb{N}$. Let $Q,D\in\mathcal{S}_{1}^{+}(H)$, $\Upsilon\in H$, and $\mathbb{A}\in\mathcal{L}(H)$. Let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup $(S_{t})_{t\geq 0}$ on $H$, and $(\mathcal{A}^{*},\operatorname{dom}(\mathcal{A}^{*}))$ its adjoint. Let $u=(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{L}_{n}^{\mathbb{C}}$ and suppose that the real Riccati system (2.14a)–(2.14c) with initial data $\Re(u)=(\Re(u_{1}),\Re(u_{2}))$ admits a (unique) mild solution $(P_{n},q_{1,n},q_{2,n})$ on $[0,T]$, as guaranteed by Proposition 5.6. Then the complex Riccati system (2.15a)–(2.15c) with initial data $u$ admits a (necessarily unique) mild solution $(\eta_{n},\zeta_{1,n},\zeta_{2,n})$ on $[0,T]$.
@@
To keep the presentation concise and highlight the numerical methodology, we approximate the Laplace transform of the infinite-dimensional Wishart stochastic volatility model directly. Let $(X_{t},\mathcal{Y}_{t})_{t\geq 0}$ be as in (2.10), and let $(\eta_{n},\zeta_{1},\zeta_{2,n})$ be the solution to (2.15) with initial condition $(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{L}_{n}^{\mathbb{C}}$, whose existence is guaranteed by Theorem 2.8.
@@
$\mathbb{E}\left[\mathop{}\!\mathrm{e}^{\langle X_{T},u_{1}\rangle_{H^{\mathbb{C}}}-\operatorname{Tr}\big(\mathcal{Y}_{T}\,\iota_{n}u_{2}\iota_{n}^{*}\big)}\,\middle|\,\mathcal{F}_{t}\right]$
@@
The approximation in (2.2.4) is the object used in the numerical section below. For general infinite-dimensional data we do not prove convergence of the right-hand side as $n\to\infty$.
@@ 3. Pricing options on forwards in infinite-dimensional stochastic volatility models
Suppose $w\colon\mathbb{R}^{+}\rightarrow[1,\infty)$ is a non-decreasing function such that $w(0)=1$ and $\int_{0}^{\infty}w^{-1}(x)\,\mathop{}\!\mathrm{d}x<\infty$. Let $H_{w}$ be the space of absolutely continuous functions $f:\mathbb{R}^{+}\rightarrow\mathbb{R}$ such that
$\|f\|_{w}^{2}:=f(0)^{2}+\int_{0}^{\infty}w(x)|f^{\prime}(x)|^{2}\,\mathop{}\!\mathrm{d}x<\infty\,$,
@@
$\langle f,g\rangle_{w}:=f(0)g(0)+\int_{0}^{\infty}w(x)f^{\prime}(x)g^{\prime}(x)\,\mathop{}\!\mathrm{d}x.$
@@
The left-shift semigroup is a strongly continuous semigroup on $H_{w}$ with infinitesimal generator $\mathcal{A}$ being the operator of differentiation (in time-to-maturity coordinate). Moreover, the point evaluation $\delta_{x}(h)=h(x)$ is a continuous linear functional on $H_{w}$, for all $x\in\mathbb{R}^{+}$. It can be expressed as (see [23, Lemma 5.3.1])
@@
where for $x\in\mathbb{R}^{+}$, $u_{x}:\mathbb{R}^{+}\rightarrow\mathbb{R}$, $y\mapsto 1+\int_{0}^{x\wedge y}w^{-1}(z)\,\mathop{}\!\mathrm{d}z$.
In particular, setting $h_{0}:=u_{0}\equiv 1\in H_{w}$, we have $u_{x}=S^{*}(x)h_{0}$ for every $x\geq 0$.
@@
$\mathop{}\!\mathrm{d}X_{t}=\big(\mathcal{A}X_{t}+C_{t}\big)\,\mathop{}\!\mathrm{d}t+D^{1/2}Y_{t}^{1/2}\,\mathop{}\!\mathrm{d}W_{t},\qquad X_{0}=x\in H_{w},$ (3.2)
where $\mathcal{A}=\partial_{x}$ is the derivative (shift) operator, $C$ is an integrable adapted $H_{w}$–valued process, $D\in\mathcal{L}(H_{w})$ is self-adjoint and positive, $(W_{t})_{t\geq 0}$ is an $H_{w}$-cylindrical Brownian motion, and $(Y_{t})_{t\geq 0}$ is either $(\mathcal{Y}_{t})_{t\geq 0}$ from (2.2.1) or $({\sf Y}_{t})_{t\geq 0}$ from (2.2). On a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{P})$, we model the (arbitrage-free) forward curve via
@@
$\mathbb{E}\left[\exp\Big\{-\int_{0}^{T}\langle\xi_{s},\mathop{}\!\mathrm{d}W_{s}\rangle_{w}-\tfrac{1}{2}\int_{0}^{T}\|\xi_{s}\|_{w}^{2}\,\mathop{}\!\mathrm{d}s\Big\}\right]=1.$ (3.4)
By Girsanov’s theorem for cylindrical Brownian motion (e.g., [18, Thm. 13]),
$\widetilde{W}_{t}:=W_{t}-\int_{0}^{t}\xi_{s}\,\mathop{}\!\mathrm{d}s$
@@
$\left.\frac{\mathop{}\!\mathrm{d}\mathbb{Q}}{\mathop{}\!\mathrm{d}\mathbb{P}}\right|_{\mathcal{F}_{t}}=\exp\Big\{-\int_{0}^{t}\langle\xi_{s},\mathop{}\!\mathrm{d}W_{s}\rangle_{w}-\tfrac{1}{2}\int_{0}^{t}\|\xi_{s}\|_{w}^{2}\,\mathop{}\!\mathrm{d}s\Big\},\qquad 0\leq t\leq T.$
@@
$\mathop{}\!\mathrm{d}X_{t}=\big(\mathcal{A}X_{t}+C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}\big)\,\mathop{}\!\mathrm{d}t+D^{1/2}Y_{t}^{1/2}\,\mathop{}\!\mathrm{d}\widetilde{W}_{t},\qquad 0\leq t\leq T.$ (3.5)
@@
Let $\xi$ satisfy (3.4). Let $\mathcal{A}$, $C$, $D$, and $(Y_{t})_{t\geq 0}$ be as above, and let $S(t)$ be the left–shift semigroup generated by $\mathcal{A}$. Recall that $u_{x}=S^{*}(x)h_{0}$ with $h_{0}\equiv 1$. Suppose that, for all $0\leq t\leq\tau\leq T$,
$C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}=-\tfrac{1}{2}\,D^{1/2}Y_{t}D^{1/2}\,u_{\tau-t}.$ (3.6)